OpenStudy (anonymous):

In a galaxy far, far away, a solar system similar to our own contains several planets. The mass of one planet in this distant galaxy is 0.7 that of the Earth's mass , and its radius is 0.51 that of the Earth's radius. 1) How much would a 78.0 kg man weigh on the surface of this far away planet? What is R, the ratio of the man's weight on this planet to his weight on Earth? 3) What is Fnew, the factor by which the man's weight is reduced if he were orbiting the planet at a distance of 3.0 Earth radii above the planet's surface?

OpenStudy (anonymous):

\[F=G(\frac{ Mm }{ R^2 })\] I am not sure of what you mean by doing a comparison

OpenStudy (anonymous):

hmm, like this example. \(x_1 y_1 =k\) --(1) \(x_2 y_2 =k\) --(2) \(\frac {(1)}{(2)}\) =>\(\frac {x_1 y_1}{x_2 y_2}=\frac{k}{k}\)

OpenStudy (anonymous):

Ok.. So I plug in the values I know for 2 seperate eqs? \[F=G \frac{ Mass Earth* Mas something }{Radius Earth ^2 }\]

OpenStudy (anonymous):

@Shadowys Is this the right approach?

OpenStudy (anonymous):

yup! :) You've got it. Though the question isn't really pretty...lol

OpenStudy (anonymous):

What would the Mass something be though? This is supposed to be a intro 100 level class. The 200 level class does easier things then this...

OpenStudy (anonymous):

@Shadowys Could use some help. I've been trying this for about 2 hrs. Still don't understand.

OpenStudy (anonymous):

@Shadowys \[F= (6.67300E-11)\frac{ 6.972E24(.70) }{6371(.58) }\]

OpenStudy (anonymous):

Forgot to mention that the top part of fraction should be multiplied by 74 kg (mass of person)

OpenStudy (anonymous):

the mass something will be cancelled out in the ratio...

OpenStudy (anonymous):

\(\frac {F_{E}}{F_{P}} =\frac {\frac{ GmM_{E}}{r^2_E}}{\frac{ GmM_{P}}{r^2_P}}\) where \(r_P = 0.51 r_E\) and \(M_P = 0.7 M_E\) also, \(F_E=mg_E\) I think this should clear things up.

OpenStudy (anonymous):

Well the answer is 2056 but I keep getting 2406

OpenStudy (anonymous):

Its ok. I got it

OpenStudy (anonymous):

I used a wrong number. The numbers were really large so I guess I put in a wrong number

OpenStudy (anonymous):

hmm wait . I got 2057N though lol

OpenStudy (anonymous):

I used 6.972 instead of the correct 5.972. The answer takes a couple digits off so its all good. Thanks

OpenStudy (anonymous):

You're welcome :)